# A Survey of Reinforcement Learning For Economics

### 09 Mar 2026 v1

@@ 11.3 The Broader Causal RL Landscape
The confounded MDP and backdoor-adjusted OPE address one specific failure mode, but the intersection of causality and reinforcement learning extends considerably further. Three active research directions illustrate the scope.
@@
Causal representation learning for RL seeks state representations $\phi(s)$ that capture causal mechanisms rather than spurious correlations. When superficial statistics change across environments but the underlying causal structure is preserved, policies learned on invariant causal features generalize where standard representations fail (Schölkopf et al., 2021). da Costa Cunha et al. (2025) formalize this as invariant policy optimization within a multi-environment MDP framework. Sim-to-real robotics illustrates the problem. A robot arm trained in a simulator with blue backgrounds and simple textures may learn to associate visual quirks with correct actions rather than learning the actual object positions that matter. When deployed on a real factory floor with different lighting, the policy fails. Training across multiple visually distinct simulators, same physics but different textures and lighting, forces the state representation to discard renderer-specific features and retain only causally relevant ones like object pose and joint angles (da Costa Cunha et al., 2025).
@@
Counterfactual policy optimization uses structural causal models to generate “what-if” trajectories for credit assignment and data augmentation. Buesing et al. (2019) introduce Gumbel-Max SCMs that produce counterfactual rollouts from a single observed trajectory, enabling off-policy evaluation without importance weights. This follows Pearl’s Abduction-Action-Prediction procedure (Pearl, 2009): first infer the exogenous noise $U$ that explains the observed trajectory (abduction), then replace the logged action with the counterfactual action (action), and finally propagate the modified action through the original structural equations to predict the counterfactual outcome (prediction). Oberst and Sontag (2019) extend this to Gumbel-Max SCM-based counterfactual off-policy estimation in healthcare settings, while Forney et al. (2017) demonstrate that counterfactual data-fusion can augment online RL exploration by generating synthetic transitions consistent with the learned causal model. In Atari Breakout, moving the paddle determines whether a brick breaks many timesteps later, but which of the dozens of intervening actions actually mattered? Buesing et al. (2019) answer this by replaying an observed trajectory with one action changed while holding the environment’s random outcomes fixed. Because the real and counterfactual trajectories share the same randomness, any difference in reward is caused by that single action change, solving the long-horizon credit assignment problem without the compounding errors of model rollouts or the high variance of importance sampling.
@@
Causal transfer in RL applies the transportability theory of Pearl and Bareinboim (2014) and Bareinboim and Pearl (2016) to policy transfer across domains. Causal invariance identifies which components of a learned policy remain valid when the environment changes, enabling transfer in settings where standard domain adaptation, which relies on distributional similarity, fails because the intervention targets differ. Autonomous driving illustrates the advantage. A self-driving policy trained in simulation has access to millions of miles of data, but the simulator’s graphics differ from real roads. Rather than treating this as one large distribution shift, selection diagrams identify which mechanisms are shared (physical dynamics of steering and braking) and which differ (visual rendering, weather effects) (Bareinboim and Pearl, 2016). The shared mechanisms transfer directly; only the differing components require recalibration with scarce real-world data.
For comprehensive surveys covering these directions and others, including fairness, safety, and partial observability, see Deng et al. (2023) and da Costa Cunha et al. (2025).
11.4 Alternative Identification Strategies
The backdoor criterion requires observing a variable $\mathbf{Z}_{t}$ that blocks all confounding paths. When no such variable is available, three alternative identification strategies apply, each with a direct econometric analogue.
When the backdoor criterion fails because confounders are unobserved, the front-door criterion offers an alternative if the action $A_{t}$ affects the next state $S_{t+1}$ only through an observed mediator $M_{t}$. The front-door criterion requires: (i) $A_{t}$ has no direct path to $S_{t+1}$ bypassing $M_{t}$, and (ii) no unblocked confounding between $M_{t}$ and $S_{t+1}$. When satisfied, a two-stage formula recovers the causal effect: first estimate how $A_{t}$ affects $M_{t}$, then how $M_{t}$ affects $S_{t+1}$ while adjusting for $A_{t}$ (Pearl, 2009). This decomposition into direct and indirect effects is the sequential analogue of mediation analysis in econometrics (da Costa Cunha et al., 2025).
The causal RL survey of da Costa Cunha et al. (2025) illustrates the front-door criterion with a mobile wellness intervention. A health app (action $A_{t}$) aims to reduce patient cortisol levels (outcome $S_{t+1}$), but unobserved health consciousness $U_{t}$ confounds the relationship because health-conscious individuals are both more likely to adopt the app and more likely to have low cortisol regardless. The app affects cortisol only through an observed mediator, supplement adherence $M_{t}$. The front-door criterion exploits this structure in two stages. First, the effect of app usage on supplement adherence is unconfounded because health consciousness does not block the causal path from $A_{t}$ to $M_{t}$. Second, the effect of supplement adherence on cortisol is identified by adjusting for app usage, which satisfies the backdoor criterion for the $M_{t}\to S_{t+1}$ link. Combining these two estimates recovers the causal effect of the app on cortisol without ever observing health consciousness.
When neither backdoor nor front-door variables are available, instrumental variable methods can identify causal effects. An instrument $Z_{t}$ must satisfy two conditions: it affects the action $A_{t}$ (relevance) and its effect on $S_{t+1}$ is channeled entirely through $A_{t}$ (exclusion restriction). Liao et al. (2024) formalize this as a Confounded MDP with Instrumental Variables (CMDP-IV), where transitions take the form $S_{t+1}=F^{*}(S_{t},A_{t})+\epsilon_{t}$ with unobserved confounders $\epsilon_{t}$ affecting both the behavior policy and transitions. Their IV-aided Value Iteration algorithm recovers the transition function $F^{*}$ from a conditional moment restriction $\mathbb{E}[F^{*}(S_{t},A_{t})\mid Z_{t}]=\mathbb{E}[S_{t+1}\mid Z_{t}]$, then runs standard value iteration on the estimated model.
Liao et al. (2024) illustrate with neonatal intensive care unit (NICU) assignment. A hospital must decide whether to admit each newborn to the NICU (action $A_{t}$), and this decision is confounded by unobserved severity indicators that affect both the admission decision and the infant’s health trajectory. Differential travel time from the hospital to a specialty care provider serves as an instrument $Z_{t}$. Relevance holds because longer travel times discourage NICU referrals, shifting the admission probability. The exclusion restriction holds because travel time affects infant health outcomes only through the admission decision, not directly. The conditional moment restriction $\mathbb{E}[F^{*}(S_{t},A_{t})\mid Z_{t}]=\mathbb{E}[S_{t+1}\mid Z_{t}]$ uses variation in travel time across hospitals to trace out the causal effect of NICU admission on health transitions, recovering the transition function $F^{*}$ that standard regression conflates with the unobserved confounder.
When confounders are truly latent but proxy variables, noisy correlates of the confounder, are available, proximal causal inference provides identification. Bennett and Kallus (2021) adapt this to sequential settings. The analyst observes two proxies: a treatment-side proxy $Z_{t}$ (e.g., previous observation $O_{t-1}$) and an outcome-side proxy $W_{t}$ (e.g., current observation $O_{t}$), both conditionally independent given the latent confounder $U_{t}$. Two bridge functions, analogous to inverse propensity scores and Q-functions, are estimated from data by solving conditional moment equations, yielding a doubly robust estimator of the policy value that is $\sqrt{n}$-consistent without ever observing $U_{t}$. Bennett and Kallus (2021) demonstrate with a sepsis management simulator in which physicians choose among fluids, vasopressors, and antibiotics at each decision point. The patient’s diabetes status is the latent confounder, partially censored in the medical record so that 20% of diabetic patients appear as non-diabetic in the data. Because the physician observes the true diabetes status when prescribing but the analyst’s dataset contains the censored version, the behavioral policy depends on a variable the analyst cannot fully recover. Previous clinical observations $O_{t-1}$ (prior lab values and vitals) serve as the treatment-side proxy $Z_{t}$ because they correlate with diabetes status and influence the physician’s current prescribing. Current clinical observations $O_{t}$ serve as the outcome-side proxy $W_{t}$ because they reflect diabetes status and predict future health transitions. In their experiments, the proximal estimator correctly identified which evaluation policy improved over the behavioral policy in 82–100% of test cases, while naive estimators that ignored confounding and standard MDP estimators that assumed full observability both achieved 0% accuracy.
11.5 Simulation Study: Confounded Off-Policy Evaluation
Unlike the simulations in Sections 56, where each algorithm generated its own training data through interaction, the estimators here receive a fixed dataset collected under the behavioral policy $\mu$. I construct a 5-state chain MDP $\mathcal{S}=\{0,1,2,3,4\}$ with two actions (advance, stay) and an absorbing goal at state 4. The causal structure at each step is $Z_{t}\to U_{t}\to A_{t}$ and $Z_{t}\to S_{t+1}$, where $Z_{t}\in\{0,1\}$ is an observed covariate drawn i.i.d. with $P(Z_{t}{=}1)=0.5$, and $U_{t}\in\{0,1\}$ is an unobserved confounder with $P(U_{t}{=}1\mid Z_{t}{=}1)=0.9$ and $P(U_{t}{=}1\mid Z_{t}{=}0)=0.1$. Transitions depend on $Z_{t}$, not $U_{t}$: the advance action succeeds with probability $0.9$ when $Z_{t}=1$ and $0.5$ when $Z_{t}=0$; the stay action is a deterministic self-loop. Rewards are a deterministic step cost r(s,a)=1r(s,a)=-1 for $s<4$, with $\gamma=0.9$. The behavioral policy depends on the unobserved $U_{t}$: $\mu(\text{advance}\mid s,U_{t})=0.55+\rho\cdot 0.25\cdot(2U_{t}-1)$, where $\rho\in\{0,0.2,\ldots,1.0\}$ controls confounding strength. The target policy always advances. Since $Z_{t}$ satisfies the backdoor criterion relative to $(A_{t},S_{t+1})$, conditioning on $Z_{t}$ blocks the confounding path $A_{t}\leftarrow U_{t}\leftarrow Z_{t}\rightarrow S_{t+1}$, and Theorem 6 guarantees identification. I compare four estimators: the oracle (analytical Bellman solution using the true interventional transition $P(s^{\prime}{=}s{+}1\mid s,\operatorname{do}(\text{advance}))=0.70$); the naive estimator (Bellman solution using observational transitions $\hat{P}(s^{\prime}\mid s,a)$ from counts, biased per Lemma L1); the backdoor-adjusted estimator ($\hat{P}(s^{\prime}\mid s,\operatorname{do}(a))=\sum_{z}\hat{P}(s^{\prime}\mid s,a,z)\,\hat{P}(z\mid s)$, which validates Theorem 6); and a naive importance-sampling estimator using marginal propensities $\hat{\mu}(a\mid s)$, biased per Lemma L1 because the true propensity $\mu(a\mid s,U_{t})$ depends on the unobserved $U_{t}$. Each configuration uses 2,000 trajectories averaged over 20 seeds.
Table 18 reports bias and root mean squared error across confounding strengths. At $\rho=0$, all model-based estimators have near-zero bias, confirming correct implementation. As $\rho$ increases, the naive estimator’s bias grows monotonically to +0.30+0.30 at $\rho=1$: the advance action is more likely when $U_{t}=1$ (high effort), which correlates with $Z_{t}=1$ (favorable conditions), so observational transitions overestimate the advance success probability ($\hat{P}_{\text{obs}}\approx 0.77$ versus the true $0.70$). This per-step bias of +0.07+0.07 compounds through the Bellman recursion; at $\rho=1$, the value bias at $s=0$ (+0.30+0.30) is nearly three times the bias at $s=3$ (+0.11+0.11), confirming the multi-period amplification that the Bellman operator produces. The backdoor-adjusted estimator eliminates this bias (RMSE <0.03<0.03 at all $\rho$), directly validating the identification formula of Theorem 6. The naive importance-sampling estimator exhibits persistent positive bias across all $\rho$ due to the mismatch between marginal and conditional propensities. Figure 9 displays the bias trajectories.
Table 18: Confounded OPE in a 5-state chain MDP: bias and RMSE of three estimators across confounding strengths $\rho$ (20 seeds, 2,000 trajectories per seed). The unobserved confounder $U_{t}$ affects the behavioral policy; the observed covariate $Z_{t}$ affects transitions and satisfies the backdoor criterion.
Naive Backdoor Naive IS
$\rho$ Bias (SE) RMSE Bias (SE) RMSE Bias (SE) RMSE
0.0 +0.006+0.006 ($0.003$) $0.014$ +0.005+0.005 ($0.003$) $0.013$ +0.635+0.635 ($0.087$) $0.740$
0.2 +0.067+0.067 ($0.004$) $0.069$ +0.003+0.003 ($0.003$) $0.015$ +0.548+0.548 ($0.117$) $0.748$
0.4 +0.125+0.125 ($0.004$) $0.127$ +0.000+0.000 ($0.004$) $0.018$ +0.654+0.654 ($0.120$) $0.837$
0.6 +0.182+0.182 ($0.003$) $0.182$ 0.004-0.004 ($0.004$) $0.017$ +0.602+0.602 ($0.093$) $0.726$
0.8 +0.242+0.242 ($0.003$) $0.242$ +0.002+0.002 ($0.004$) $0.019$ +0.593+0.593 ($0.085$) $0.699$
1.0 +0.300+0.300 ($0.004$) $0.301$ +0.008+0.008 ($0.006$) $0.026$ +0.589+0.589 ($0.094$) $0.718$
Figure 9: Bias of three OPE estimators relative to the oracle as a function of confounding strength $\rho$. Naive bias grows monotonically as the behavioral policy’s dependence on the unobserved confounder $U_{t}$ increasingly distorts observational transition estimates. The backdoor-adjusted estimator (Theorem 6) remains unbiased at all confounding levels.
@@ 12 Conclusion
Economics has long relied on dynamic programming to formalize and solve problems of sequential decision-making. Yet, for decades, the complexity of our models has been dictated less by the economic phenomena we wish to study and more by the computational limits of the curse of dimensionality. Reinforcement learning offers one pragmatic path forward. By replacing exhaustive state-space enumeration with sampling and function approximation, RL trades the absolute convergence guarantees of classical dynamic programming for the capacity to operate in high-dimensional, continuous, and complex strategic environments.

### 15 Mar 2026 v2

@@ 13.5 The Broader Causal RL Landscape
Three active research directions beyond confounded MDPs illustrate the broader scope of causal RL.
@@
Causal representation learning for RL seeks state representations $\phi(s)$ that capture causal mechanisms rather than spurious correlations (Schölkopf et al., 2021). da Costa Cunha et al. (2025) formalize this as invariant policy optimization within a multi-environment MDP framework. In sim-to-real robotics, training across multiple visually distinct simulators (same physics, different rendering) forces the representation to discard renderer-specific features and retain only causally relevant ones like object pose and joint angles.
@@
Counterfactual policy optimization uses structural causal models to generate “what-if” trajectories for credit assignment and data augmentation. Buesing et al. (2019) introduce Gumbel-Max SCMs that produce counterfactual rollouts from a single observed trajectory via Pearl’s Abduction-Action-Prediction procedure (Pearl, 2009): infer the exogenous noise explaining the observed trajectory, replace the logged action, and propagate through the structural equations. By replaying a trajectory with one action changed while holding the environment’s randomness fixed, any difference in reward is attributable to that action, solving long-horizon credit assignment without importance sampling. Oberst and Sontag (2019) extend this to healthcare settings, while Forney et al. (2017) use counterfactual data-fusion to augment online exploration.
@@
Causal transfer in RL applies the transportability theory of Pearl and Bareinboim (2014) and Bareinboim and Pearl (2016) to policy transfer across domains. Selection diagrams identify which mechanisms are shared across environments and which differ, enabling targeted recalibration rather than wholesale domain adaptation. In sim-to-real autonomous driving, physical dynamics transfer directly while visual rendering requires recalibration with scarce real-world data (Bareinboim and Pearl, 2016).
13.6 Simulation Study: Confounded Retail Pricing MDP
I construct a 5-state engagement funnel $\mathcal{S}=\{0,1,2,3,4\}$ with two actions (promote, hold price) and an absorbing conversion state at $s=4$. A retailer manages customers through engagement stages, deciding whether to offer a promotional discount. The data-generating process embeds four distinct sources of causal variation, enabling simultaneous validation of all four identification strategies from a single DGP.
[…]
Figure 26: Complete causal graph of the simulation DGP. Gray dashed node ($U_{t}$) is unobserved; dashed edges involve $U_{t}$.
The action affects the next state only through the mediator: $P(s{+}1\mid s,M_{t},Z_{t})$ depends on $M_{t}$ and $Z_{t}$ but not on $A_{t}$ or $U_{t}$ directly. This enables all four identification strategies simultaneously: $Z_{t}$ satisfies the backdoor criterion, $M_{t}$ satisfies the front-door criterion, $\text{IV}_{t}$ satisfies relevance and exclusion, and $W_{t}^{(1)},W_{t}^{(2)}$ satisfy the proximal conditions.207207207Rewards are r(s,a)=1r(s,a)=-1 for $s<4$, $\gamma=0.9$. The target policy always promotes. The true interventional transition probability is $P(s{+}1\mid s,\operatorname{do}(\text{promote}))=0.615$.
I compare six estimators: oracle, naive (biased per Lemma L1), backdoor (Equation 111), front-door (Equation 112), Wald IV (Equation 114), and proximal (Equations 115116).208208208Each configuration uses 2,000 trajectories averaged over 20 seeds.
Table 26 reports bias and RMSE across confounding strengths. The naive estimator’s bias grows monotonically with $\rho$ because observational transitions overestimate promotion success: the promote action is more likely when $U_{t}=1$, which correlates with favorable conditions, so $\hat{P}_{\text{obs}}$ exceeds the true interventional probability, and this per-step bias compounds through the Bellman recursion. The backdoor and front-door estimators eliminate bias at all $\rho$, validating Theorem 8 and Equation (112). The IV estimator maintains low bias but higher variance due to the Wald ratio’s sensitivity to instrument strength; panel (b) of Figure 27 illustrates this classic bias-variance tradeoff. The proximal estimator achieves low bias with moderate variance, confirming that bridge functions recover causal effects from noisy proxies.
Table 26: Unified identification in a confounded retail pricing MDP: bias and RMSE of five estimators across confounding strengths $\rho$ (20 seeds, 2,000 trajectories per seed).
Naive Backdoor Front-door IV Proximal
$\rho$ Bias (SE) RMSE Bias (SE) RMSE Bias (SE) RMSE Bias (SE) RMSE Bias (SE) RMSE
0.0 0.010-0.010 ($0.005$) $0.024$ 0.008-0.008 ($0.005$) $0.024$ 0.001-0.001 ($0.005$) $0.023$ 0.006-0.006 ($0.028$) $0.122$ 0.011-0.011 ($0.006$) $0.028$
0.2 +0.060+0.060 ($0.005$) $0.064$ 0.007-0.007 ($0.006$) $0.025$ +0.003+0.003 ($0.006$) $0.025$ +0.012+0.012 ($0.028$) $0.125$ 0.012-0.012 ($0.006$) $0.027$
0.4 +0.127+0.127 ($0.005$) $0.129$ 0.008-0.008 ($0.006$) $0.026$ +0.001+0.001 ($0.006$) $0.027$ +0.011+0.011 ($0.029$) $0.127$ 0.015-0.015 ($0.006$) $0.028$
0.6 +0.194+0.194 ($0.005$) $0.195$ 0.007-0.007 ($0.006$) $0.028$ +0.001+0.001 ($0.006$) $0.028$ +0.020+0.020 ($0.030$) $0.131$ 0.014-0.014 ($0.007$) $0.033$
0.8 +0.258+0.258 ($0.005$) $0.259$ 0.007-0.007 ($0.006$) $0.029$ 0.004-0.004 ($0.007$) $0.030$ +0.032+0.032 ($0.030$) $0.136$ 0.015-0.015 ($0.007$) $0.034$
1.0 +0.321+0.321 ($0.005$) $0.322$ 0.005-0.005 ($0.006$) $0.028$ 0.004-0.004 ($0.007$) $0.030$ +0.055+0.055 ($0.025$) $0.123$ 0.015-0.015 ($0.008$) $0.037$
Figure 27: (a) Bias of five OPE estimators relative to the oracle as a function of confounding strength $\rho$. (b) IV estimator bias distribution vs. instrument strength at $\rho=1$. Dashed red line: naive estimator bias.
@@ 14 Conclusion
Economics has long relied on dynamic programming to formalize and solve problems of sequential decision-making. Yet, for decades, the complexity of our models has been dictated less by the economic phenomena we wish to study and more by the computational limits of the curse of dimensionality.209209209Rust (2019) asks the same question of dynamic programming itself, surveying its practical impact across domains and identifying the gap between theoretical power and real-world deployment. Reinforcement learning offers one pragmatic path forward. By replacing exhaustive state-space enumeration with sampling and function approximation, RL trades the absolute convergence guarantees of classical dynamic programming for the capacity to operate in high-dimensional, continuous, and complex strategic environments.